# Surface topological quantum criticality: Conformal manifolds and Discrete Strong Coupling Fixed Points

### v1

@@ I Introduction
In bulk lattices, SUSY can also emerge with two half-Dirac cones in a bulk honeycomb lattice[48]. The kinematic constraint in the lattice effectively forbids the coupling between two half-Dirac cones resulting in two separated copies of SUSY theories.
@@
In this article, we will focus on more general topological surface states that can appear in a topological insulator with a crystal translation symmetry where the number of Dirac cones is defined as $N_{f}=\frac{\mathcal{N}}{2}$, with $\mathcal{N}=1,2,3,4$. We will exclusively focus on the cases of $\mathcal{N}=2,3$ where conformal manifolds of strong coupling fixed points naturally appear on phase boundaries of gapless free fermions. All the CFTs we discuss in this article do not exhibit an SUSY, unlike in the case of $\mathcal{N}=1$; however, they represent generic strongly interacting topological surfaces of a topological cubic crystal.
@@
Section III is dedicated to the exploration of various topological phases of a simple 3D cubic lattice model and their surface states. Being the surface of a 3D topological insulator, it could host an $\mathcal{N}$ number of surface half-Dirac cones (from now on we will use $\mathcal{N}=2N_{f}$ to count the degrees of surface fermions) where $\mathcal{N}$ can be $1$, $2$ or $3$. For the trivial phase, $\mathcal{N}$ can be either 00 or $4$. Thus we focus only on the cases of $\mathcal{N}$ being $1,2,3$. If $\mathcal{N}$ is odd(even), then the material is a strong(weak) TI. It turns out that the topological classification does not uniquely identify the exact number of surface Dirac cones, rather it defines the even-odd parity of $\mathcal{N}$. Hence, it is possible that for the same bulk topological invariant, the surface can host different numbers of Dirac fermions, as long as the count remains odd or even. In this context, the stability of the number $\mathcal{N}$ of the surface Dirac fermions comes into question. This section addresses the issue, demonstrating that the exact number of surface Dirac cones (denoted by $\mathcal{N}$) can be stabilized by imposing additional symmetry constraints associated with a crystal (crystal translation symmetry, inversion symmetry, etc), despite identical bulk topological $Z_{2}$ invariants but different surfaces. This is very crucial and significant for our later discussions on phase boundaries and conformal manifolds of fixed points that define the universality of phase transitions across the phase boundaries. As we shall show in this article, the $\mathcal{N}=1$ and $\mathcal{N}=3$ surfaces can have different phase boundary manifolds, placing them in distinctly different universality classes.
@@
Table 1: Summary of results: List of isolated fixed points, conformal manifolds and phase boundary of an attractively interacting TI (cubic crystal) surface with $\mathcal{N}$ Dirac cones when studied upto one-loop order. Here $\hat{V}^{\mathcal{N}\times\mathcal{N}}_{c}$ is the critical interaction matrix of size $\mathcal{N}$ (see Eqn.27). $\textbf{n}_{\mathcal{N}}$ is a unit vector in $\mathcal{N}$-dimensional space.
@@
In section IV, we introduce effective attractive interactions to the 3D TI with an arbitrary $\mathcal{N}$ number of surface Dirac cones. We first develop an effective four-fermion field theory in $2+1D$ for the TI surface with $\mathcal{N}$ 2-component Dirac fermions and then use the renormalization group equation (RGE) approach in $2+\epsilon$ dimensions to study the surface quantum criticality. We find that the RG equation possesses an emergent $SO(\mathcal{N})$ symmetry at the $\mathcal{O}(\epsilon)$ order or the one-loop approximation while it is broken at the $\mathcal{O}(\epsilon^{2})$ order or the two-loop approximation. In the spirit of $\epsilon$-expansion, in this section we limit our studies to the one-loop approximation, as a consequence of which the RGE possesses an emergent $SO(\mathcal{N})$ symmetry.
@@
This emergent symmetry of the dynamical equation results in certain classes of fixed points forming a conformal manifold in the parameter space, for $\mathcal{N}>1$. The geometry of the manifold is determined by the coset space formed by dividing the symmetry group of the RG equation ($SO(\mathcal{N})$ group in our case) with the invariant subgroup of the scale symmetric solutions. Section V is dedicated to the discussion of this phenomenon in detail.
@@
As stated before, the $\mathcal{N}=1$ limit is well studied and it is found that the surface possesses a UV fixed point in addition to the free-fermion fixed point. In section VI, we discuss the fixed points of the theory when $\mathcal{N}=2$, that is when there are two flavors of surface fermions. We observe that in addition to the free-fermion and the unstable strong-coupling fixed points, there exists a manifold of interacting fixed points that forms a ring in the parameter space. This ring is embedded in a 2-dimensional conical phase boundary separating the gapless and superconducting phases. The infrared physics of the conical phase boundary can be described effectively by a Yukawa field theory which involves two flavors of surface fermions coupling with a single flavor of complex bosons.
@@
In section VII, we discuss the types of fixed points when there are three flavors of surface fermions on the TI surface. We observe that there are two conformal manifolds both in the shape of a 2-sphere. The 2-sphere conformal manifolds are embedded in a 5-dimensional phase boundary. Similar to the $\mathcal{N}=2$ case, the effective Yukawa field theory of the 5-dimensional phase boundary involves $three$ flavors of surface fermions coupling with a single flavor of complex bosons.
@@
In our current studies, we also find that the emergent symmetry in the RGEs, though a sufficient condition, is not a necessary condition for the formation of conformal manifolds. We show this explicitly in section VIII, where certain three-loop effects are studied for the $\mathcal{N}=2$ TI surface. A smooth manifold persists in the presence of such $SO(\mathcal{N})$ symmetry-breaking effects, albeit in a distorted shape. However, in general, the conformal manifolds can further break down to a few isolated fixed points when $SO(\mathcal{N})$ symmetry-breaking effects are included.
In section IX, we conclude our studies and discuss future directions.
@@ II Summary of main results
The phase boundaries of the gapless weakly interacting topological fermions strongly depend on the number of surface Dirac cones, $\mathcal{N}$. The dimension of the interaction parameter space (indicated by the subscript $p$ below) in our studies is $D_{p}=\mathcal{N}(\mathcal{N}+1)/2$ (see section VI,VII). $D_{p}=3$ and $6$ respectively when $\mathcal{N}=2,3$ which we will focus on. The generic phase boundaries are given by manifolds of co-dimension $D_{co}=1$ embedded in a $D_{p}$ dimension parameter space.
@@
All the quantum phase transitions represented by any points in the phase boundary manifold are characterized by these conformal manifolds identified in our study. See details of the operator scaling dimensions near conformal manifolds (Appendix.C).
@@
III 3D Topological insulators: Distinct surface states and deformation between them
@@
Figure 1: a) Phase diagram of the 3D topological insulator (TI) defined by the Hamiltonian in Eqn.1, with $t_{x}=t_{y}=t$. Each topological phase is distinguished by its color, and the topological invariant is denoted by $(v_{0};v_{1},v_{2},v_{3})$. b) The surface spectrum of a bulk topological phase characterized by $(1;0,0,0)$ (represented by the blue-colored region) when $t<0$. Three surface Dirac cones are observed. c) The surface spectrum for the same bulk topological phase $(1;0,0,0)$ when $t>0$. Only one single surface Dirac cone is observed. The inset illustrates the surface Brillouin zone (BZ) and its time-reversal invariant (TRI) points $\textbf{Q}_{i}(i=1,2,3,4)$.
@@
Consider the general case of a TI surface with $\mathcal{N}$ Dirac cones where $\mathcal{N}\leq 4$. We ask how robust and stable the number $\mathcal{N}$ of the surface Dirac fermions are, and what is its relation to the symmetry group of the bulk? We shall answer this question in this section.
First, using a simple bulk lattice model, we show explicitly that a strong TI can host single or three Dirac cones for the same bulk topological invariant (section.IIIA). Then, we shall show that the two phases can be connected smoothly by a unitary transformation. However, this transformation comes at the cost of breaking translation and inversion symmetry and introducing long-range hopping (see Fig.4) (section.IIIA). Consequently, we argue that introducing crystal symmetry constraints can stabilize the number of surface fermions $\mathcal{N}$ on the TI surface.
III.1 Distinct surface states for same bulk topological invariant: a practical example
@@
$\mathcal{H}(\textbf{k})$ $=$ $\sum_{i=x,y,z}\alpha_{i}\sin k_{i}+\beta\left[1-\sum_{i=x,y,z}t_{i}\cos k_{i}\right]$
@@
III.2 Distinct surface states: A consequence of flipping of mass signs
Here, we study the bulk topological invariants and examine the underlying physics that leads to distinct surface states for the same topological bulk. We demonstrate that the transition between the distinct surface states is made possible by flipping the fermion mass signs $\delta_{\Gamma_{i}}$’s (defined in Eqns.2a,2b) at the TRI points on any one of the three $k_{a}=\pi(a=x,y,z)$ planes in bulk BZ (see Fig.2).
@@ III.2 Distinct surface states: A consequence of flipping of mass signs
It is clear from Eqn.2a that the system is topologically non-trivial for an odd number of negative fermion mass signs $\delta_{\Gamma_{i}}$’s; hence, a gapless surface state exists. The other three invariants ($v_{1},v_{2},v_{3}$) correspond to the product of $\delta_{\Gamma_{i}}$’s at the four TR-invariant points on the three distinct $k_{a}=\pi$ planes (here $a=x,y,z$ respectively for $v_{1},v_{2},v_{3}$).
Figure 2: The figure illustrates how flipping the mass signs on one of the planes in the bulk Brillouin zone (BZ) leads to a change in the number of surface Dirac cones without altering the bulk topological invariant. a) Mass signs ($\delta_{\Gamma_{i}}$) at the eight TRI points in the bulk BZ. The TI phase is $(1;1,1,1)$. b) Mass signs ($\pi_{\Lambda_{i}}$) at the four TRI momenta in the surface Brillouin zone in the xy-plane. c) Flipping of mass signs at the $k_{z}=\pi$ plane.for the same bulk topological phase $(1;1,1,1)$. d) Resultant mass signs ($\pi_{\Lambda_{i}}$) in the xy-plane after flipping the mass signs. (See Eqns.2a,2b,3 for details).
@@
We use this procedure to locate the gapless points on the surface BZ in Fig. 2. In Fig. 2(a), there is a single negative sign at the $(\pi,\pi,\pi)$ point in the BZ. The bulk topological invariant is $(1;1,1,1)$. Using Eqn.3, we find that the TR polarization value $\pi_{\Lambda_{i}}=-1$ at the surface TRI point $\Lambda_{i}=(\pi,\pi)$ as shown in Fig.2(b). As a result, there must be a Dirac cone at the $(\pi,\pi)$ point in the surface BZ.
@@
III.3 Unitary deformation between distinct surface states
Here, we shall develop an explicit unitary operator that can smoothly connect the two distinct surface states of a TI with the same bulk topological invariant. We consider the 3D TI to have inversion symmetry so that $\delta_{\Gamma_{i}}$ (see Eqn.2a) is just the sign of the mass gap (parity eigenvalue) at the TRI point $\Gamma_{i=(n_{1},n_{2},n_{3})}$.
III.3.1 Effective field theory approach

### v3

@@ I Introduction
It has been known that SPT surfaces can be identified via the method of gauge anomalies (or vice versa, see Ref.[40]) as the lattice symmetry can be implemented in a non-local way on surfaces. For given anomalies, one can identify phases of surfaces depending on whether spontaneous symmetry breaking(SSB) or topological orders (TO) further appear on a surface. The surfaces of 3D bosonic SPT states and quantum criticality and dynamics have been thoroughly studied in Ref.[41]. We refer the readers to those discussions.
The current studies are fully consistent with this general paradigm of topological surfaces. What we have addressed here is that when interactions are attractive pairing ones, the classes of CFTs that appear on topological surfaces can be further directly related to bulk invariants, which we have used to infer the number of surface fermions. For instance, for strong TIs (with cubic crystal symmetry), the CFTs are either related to an isolated SUSY fixed point or a conformal manifold with its tangent space set by two marginal operators but without SUSY. These are the general results within a given class of interacting topological surfaces rather than within a given broader class of anomalies. Our family of gapless states are therefore a subclass with further inputs of interactions. They are related to quantum phase transitions to SSB states that break the protecting charge $U(1)$ symmetries.
There can also be other fascinating transitions, such as SSB to SSB states and TO to TO states. They are very interesting to look further into in connection to conformal manifolds. In our current studies, we have restricted ourselves to quantum phase transitions between the well-observed gapless free fermions phase and SSB states breaking the charge $U(1)$ symmetries with interaction-parameter inputs, in addition to topological data. Our main focus is to understand the phase boundary that separates the gapless phase and SSB states, breaking one of the protecting symmetries and what we can have on the phase boundary manifold.
@@
In lattices, SUSY can also emerge with two half-Dirac cones in a bulk honeycomb lattice[50]. The kinematic constraint in the lattice effectively forbids the coupling between two half-Dirac cones resulting in two separated copies of SUSY theories.
@@
In this article, we will focus on more general topological surface states that can appear in a topological insulator with a crystal translation symmetry where the number of Dirac cones is defined as $N_{f}=\frac{\mathcal{N}}{2}$, with $\mathcal{N}=1,2,3,4$. We will exclusively focus on the cases of $\mathcal{N}=2,3$ where conformal manifolds of strong coupling fixed points naturally appear on phase boundaries of gapless free fermions. All the CFTs we discuss in this article do not exhibit SUSY, unlike in the case of $\mathcal{N}=1$; however, they represent generic strongly interacting topological surfaces of a topological cubic crystal.
We have dedicated section.III exclusively for the discussions on the conformal manifolds. There, we discuss certain features of the conformal manifold that appear at the phase boundary of the gapless surface fermions. We focus on two aspects: 1) relations to generic conformal manifolds discussed in quantum field theories and 2) its distinction from conventional isolated fixed points.
@@
Section IV is dedicated to the exploration of various topological phases of a simple 3d cubic lattice model and their surface states. Being the surface of a 3d topological insulator, it could host an $\mathcal{N}$ number of surface half-Dirac cones (from now on we will use $\mathcal{N}=2N_{f}$ to count the degrees of surface fermions) where $\mathcal{N}$ can be $1$, $2$ or $3$. For the trivial phase, $\mathcal{N}$ can be either 00 or $4$. Thus we focus only on the cases of $\mathcal{N}$ being $1,2,3$. If $\mathcal{N}$ is odd(even), then the material is a strong(weak) TI. It turns out that the topological classification does not uniquely identify the exact number of surface Dirac cones, rather it defines the even-odd parity of $\mathcal{N}$. Hence, it is possible that for the same bulk topological invariant, the surface can host different numbers of Dirac fermions, as long as the count remains odd or even. In this context, the stability of the number $\mathcal{N}$ of the surface Dirac fermions comes into question. This section addresses the issue, demonstrating that the exact number of surface Dirac cones (denoted by $\mathcal{N}$) can be stabilized by imposing additional symmetry constraints associated with a crystal (crystal translation symmetry, inversion symmetry, etc), despite identical bulk topological $Z_{2}$ invariants but different surfaces. This is very crucial and significant for our later discussions on phase boundaries and conformal manifolds of fixed points that define the universality of phase transitions across the phase boundaries. As we shall show in this article, the $\mathcal{N}=1$ and $\mathcal{N}=3$ surfaces can have different phase boundary manifolds, placing them in distinctly different universality classes.
@@
Table 1: Summary of results: List of isolated fixed points, conformal manifolds and phase boundary of an attractively interacting TI (cubic crystal) surface with $\mathcal{N}$ Dirac cones when studied upto one-loop order. Here $\hat{V}^{\mathcal{N}\times\mathcal{N}}_{c}$ is the critical interaction matrix of size $\mathcal{N}$ (see Eqn.28). $\textbf{n}_{\mathcal{N}}$ is a unit vector in $\mathcal{N}$-dimensional space.
@@
In section V, we introduce effective attractive interactions to the 3D TI with an arbitrary $\mathcal{N}$ number of surface Dirac cones. We first develop an effective four-fermion field theory in $2+1D$ for the TI surface with $\mathcal{N}$ 2-component Dirac fermions and then use the renormalization group equation (RGE) approach in $2+\epsilon$ dimensions to study the surface quantum criticality. We find that the RG equation possesses an emergent $SO(\mathcal{N})$ symmetry at the $\mathcal{O}(\epsilon)$ order or the one-loop approximation while it is broken at the $\mathcal{O}(\epsilon^{2})$ order or the two-loop approximation. In the spirit of $\epsilon$-expansion, in this section we limit our studies to the one-loop approximation, as a consequence of which the RGE possesses an emergent $SO(\mathcal{N})$ symmetry.
@@
This emergent symmetry of the dynamical equation results in certain classes of fixed points forming a conformal manifold in the parameter space, for $\mathcal{N}>1$. The geometry of the manifold is determined by the coset space formed by dividing the symmetry group of the RG equation ($SO(\mathcal{N})$ group in our case) with the invariant subgroup of the scale symmetric solutions. Section VI is dedicated to the discussion of this phenomenon in detail.
@@
As stated before, the $\mathcal{N}=1$ limit is well studied and it is found that the surface possesses a UV fixed point in addition to the free-fermion fixed point. In section VII, we discuss the fixed points of the theory when $\mathcal{N}=2$, that is when there are two flavors of surface fermions. We observe that in addition to the free-fermion and the unstable strong-coupling fixed points, there exists a manifold of interacting fixed points that forms a ring in the parameter space. This ring is embedded in a 2-dimensional conical phase boundary separating the gapless and superconducting phases. The infrared physics of the conical phase boundary can be described effectively by a Yukawa field theory which involves two flavors of surface fermions coupling with a single flavor of complex bosons.
@@
In section VIII, we discuss the types of fixed points when there are three flavors of surface fermions on the TI surface. We observe that there are two conformal manifolds both in the shape of a 2-sphere. The 2-sphere conformal manifolds are embedded in a 5-dimensional phase boundary. Similar to the $\mathcal{N}=2$ case, the effective Yukawa field theory of the 5-dimensional phase boundary involves $three$ flavors of surface fermions coupling with a single flavor of complex bosons.
@@
In our current studies, we also find that the emergent symmetry in the RGEs, though a sufficient condition, is not a necessary condition for the formation of conformal manifolds. We show this explicitly in section IX, where certain three-loop effects are studied for the $\mathcal{N}=2$ TI surface. A smooth manifold persists in the presence of such $SO(\mathcal{N})$ symmetry-breaking effects, albeit in a distorted shape. However, in general, the conformal manifolds can further break down to a few isolated fixed points when $SO(\mathcal{N})$ symmetry-breaking effects are included.
In section X, we conclude our studies and discuss future directions.
@@ II Summary of main results
The phase boundaries of the gapless weakly interacting topological fermions strongly depend on the number of surface Dirac cones, $\mathcal{N}$. The dimension of the interaction parameter space (indicated by the subscript $p$ below) in our studies is $D_{p}=\mathcal{N}(\mathcal{N}+1)/2$ (see section VII,VIII). $D_{p}=3$ and $6$ respectively when $\mathcal{N}=2,3$ which we will focus on. The generic phase boundaries are given by manifolds of co-dimension $D_{co}=1$ embedded in a $D_{p}$ dimension parameter space.
@@
All the quantum phase transitions represented by any points in the phase boundary manifold are characterized by these conformal manifolds identified in our study. See details of the operator scaling dimensions near conformal manifolds (Appendix.D).
@@
III Conformal manifolds
Previously, we mentioned that the phase boundary manifold of the $\mathcal{N}=2,3$ flavor TI surfaces contains conformal sub-manifolds with co-dimensions $D_{co}=1,4$ respectively. These conformal manifolds observed at the surface quantum criticality require special attention. In this section, we shall describe certain generic features of the conformal manifold, especially on how it is distinct from a conventional isolated fixed point. Following this, we list down the special properties of the conformal manifold observed in our studies.
A conformal manifold is a family of interacting Hamiltonians connected by exactly marginal operators. For instance, let $\mathcal{S}_{c}$ be an action defined at a point on the conformal manifold. Let $\{\mathcal{O}^{i}_{M}\}(i=1,..n)$ be the exactly marginal operators at the given point on the manifold. Then, after a marginal deformation,
$\mathcal{S}_{c}+\delta\mathcal{S}=\mathcal{S}_{c}+\sum^{n}_{i=1}\int d^{D}x\,\delta\lambda^{i}_{M}\mathcal{O}^{i}_{M}(x)$ (1)
the theory remains conformal invariant. Here, $\delta\lambda^{i}_{M}$ is the infinitesimal marginal coupling away from the fixed point corresponding to the marginal operator $\mathcal{O}^{i}_{M}$. The number of marginal operators determines the dimension of the conformal manifold. One important property of these exactly marginal operators is that their scaling dimension $\triangle_{\mathcal{O}_{M}}=D$, the spacetime dimensions of the theory. In contrast, the canonical scaling dimensions of $\mathcal{O}_{M}$ is $22D-$ 2.
Eqns.37 and 81 provide the detailed expressions for the marginal operators $\mathcal{O}^{i}_{M}(x)$ of the conformal manifolds that appears on the phase boundary of $\mathcal{N}=2$ and $\mathcal{N}=3$ TI surfaces respectively.
Before proceeding further and starting our discussions on our results, we want to remark on the nature of the specific conformal manifolds discussed in this article and their relations to more generic conformal manifolds, which are well discussed in modern quantum field theories. As discussed above, typical construction of conformal manifolds usually starts with a set of exact marginal operators that form a tangent space of a manifold and can then be applied to generate other distinct conformal field theories via deformations.
Practical implementations of such procedures either require primary conformal operators with a known scale dimension matching the space-time dimension as in $(1+1)D$, or supersymmetry conformal operators to ensure such exact marginal operators as in a higher dimension. The best well-known high-dimension example is the N=4 supersymmetric Yang-Mills theory[55, 56], which can support un-compact conformal manifolds with logarithm divergent diameters measured using the standard Zamodolchikov metric[57].
The tangent space of our manifolds can be defined by marginal operators discussed in Appendix D when they become exactly marginal in the limit of our interests. Our manifolds can then be constructed in a similar way via deformations. However, in this case, deformations simply lead to theories within the same universality class because of the higher symmetry of our manifolds. The manifold is either a one-sphere with a uniform metric or a two-sphere with a uniform curvature tensor, i.e. they form representations of $SO(2)$ and $SO(3)$ groups, respectively. The deformation of original theories induces a rotation in the interaction parameter space. Because of the emergent $SO(\mathcal{N})$ dynamic symmetry in RGEs discussed below, the operator scaling dimensions are invariant under rotations, hence the deformations. Therefore, the deformations, although resulting in a family of distinct interaction Hamiltonians, do not result in a new universality class.
The conformal manifold in this article only hosts a single universality class, far fewer classes than the generic ones discussed before. For this reason, we name it a featureless conformal manifold to distinguish it from more generic cases, such as the ones in SCFTs, where each point on a manifold actually defines a distinct class of CFTs.
It shall be emphasized here that our result is distinctly different from an isolated Wilson-Fisher fixed point, which can be considered a zero-sphere. An isolated fixed point forbids exact marginal operators as otherwise, it again can be deformed and induce a smooth manifold. So, from this point of view, the key distinction between a smooth manifold, as discussed in our case, and an isolated point, a more standard critical point, is the existence of a tangent space at each point of the manifold and exact marginal operators, which usually define the metric of the manifold. And the number of exact marginal operators defines the dimension of the manifold just like in all other general discussions.
IV 3D Topological insulators: Distinct surface states and deformation between them
@@ III Conformal manifolds
Figure 1: a) Phase diagram of the 3d topological insulator (TI) defined by the Hamiltonian in Eqn.2, with $t_{x}=t_{y}=t$. Each topological phase is distinguished by its color, and the topological invariant is denoted by $(v_{0};v_{1},v_{2},v_{3})$. b) The surface spectrum of a bulk topological phase characterized by $(1;0,0,0)$ (represented by the blue-colored region) when $t<0$. Three surface Dirac cones are observed. c) The surface spectrum for the same bulk topological phase $(1;0,0,0)$ when $t>0$. Only one single surface Dirac cone is observed. The inset illustrates the surface Brillouin zone (BZ) and its time-reversal invariant (TRI) points $\textbf{Q}_{i}(i=1,2,3,4)$.
@@
First, using a simple bulk lattice model, we show explicitly that a strong TI can host single or three Dirac cones for the same bulk topological invariant (section.IVA). Then, we shall look for unitary operations that could connect these distinct surface states. We find two non-local unitary operators that are equivalent to infinite-depth ones. The first one (Eqn.18a) does change the bulk invariant, as anticipated, therefore connecting two topologically different TIs, which is not surprising given the general paradigm about topological bulks. The second one (Eqn.18c), however, does not change the bulk, though of infinite depth, but changes the surface Dirac cones.
The main purpose of this part of the discussion is to understand how stable the surface Dirac cones are, given that for a given bulk, there are two different surfaces. Our analysis shows that although they are associated with the same bulk, for a given topological bulk, there appears to be no other way to deform one surface into the other surface unless it is unitary with infinite depth. Here, we have constructed the second type of infinite depth unitary that changes the surface of Dirac cones without changing the bulk.
Specifically, we find that our unitary accomplishing this necessarily partially breaks protecting symmetries and induces long-range hopping. This strongly implies that when restricted to local unitary when all protecting symmetries are enforced, the surface states remain distinct and topologically stable, reinforcing our focus on the criticality of a fixed number $\mathcal{N}$ of Dirac fermions. This is further illustrated in Fig.4.
IV.1 Distinct surface states for same bulk topological invariant: a practical example
@@
$\mathcal{H}_{0}(\textbf{k})$ $=$ $\sum_{i=x,y,z}\alpha_{i}\sin k_{i}+\beta\left[1-\sum_{i=x,y,z}t_{i}\cos k_{i}\right]$
@@
IV.2 Distinct surface states: A consequence of flipping of mass signs
Here, we study the bulk topological invariants and examine the underlying physics that leads to distinct surface states for the same topological bulk. We demonstrate that the transition between the distinct surface states is made possible by flipping the fermion mass signs $\delta_{\Gamma_{i}}$’s (defined in Eqns.3a,3b) at the TRI points on any one of the three $k_{a}=\pi(a=x,y,z)$ planes in bulk BZ (see Fig.2).
@@ IV.2 Distinct surface states: A consequence of flipping of mass signs
It is clear from Eqn.3a that the system is topologically non-trivial for an odd number of negative fermion mass signs $\delta_{\Gamma_{i}}$’s; hence, a gapless surface state exists. The other three invariants ($v_{1},v_{2},v_{3}$) correspond to the product of $\delta_{\Gamma_{i}}$’s at the four TR-invariant points on the three distinct $k_{a}=\pi$ planes (here $a=x,y,z$ respectively for $v_{1},v_{2},v_{3}$).
Figure 2: The figure illustrates how flipping the mass signs on one of the planes in the bulk Brillouin zone (BZ) leads to a change in the number of surface Dirac cones without altering the bulk topological invariant. a) Mass signs ($\delta_{\Gamma_{i}}$) at the eight TRI points in the bulk BZ. The TI phase is $(1;1,1,1)$. b) Mass signs ($\pi_{\Lambda_{i}}$) at the four TRI momenta in the surface Brillouin zone in the xy-plane. c) Flipping of mass signs at the $k_{z}=\pi$ plane for the same bulk topological phase $(1;1,1,1)$. d) Resultant mass signs ($\pi_{\Lambda_{i}}$) in the xy-plane after flipping the mass signs. (See Eqns.3a,3b,4 for details).
@@
We use this procedure to locate the gapless points on the surface BZ in Fig. 2. In Fig. 2(a), there is a single negative sign at the $(\pi,\pi,\pi)$ point in the BZ. The bulk topological invariant is $(1;1,1,1)$. Using Eqn.4, we find that the TR polarization value $\pi_{\Lambda_{i}}=-1$ at the surface TRI point $\Lambda_{i}=(\pi,\pi)$ as shown in Fig.2(b). As a result, there must be a Dirac cone at the $(\pi,\pi)$ point in the surface BZ.
@@
IV.3 Unitary deformation between distinct surface states
Here, we shall develop an explicit unitary operator that can smoothly connect the two distinct surface states of a TI with the same bulk topological invariant. We consider the 3D TI to have inversion symmetry so that $\delta_{\Gamma_{i}}$ (see Eqn.3a) is just the sign of the mass gap (parity eigenvalue) at the TRI point $\Gamma_{i=(n_{1},n_{2},n_{3})}$.
IV.3.1 Effective field theory approach